OpenStudy (anonymous):

Bonnie has one alloy containing 23% aluminum and another containing 45% aluminum. How many pounds of each alloy must he use to make 58 pounds of a third alloy containing 42% aluminum?

OpenStudy (anonymous):

You can work this out using a system of equations. x represents the pounds of alloy 1 y represents the pounds of alloy 2 x + y = 58 (the number of pounds of each allow to make 58 pounds) 0.23x + 0.45y = 0.42 x 58 (the amount of aluminum in the first, second and resulting alloy.

OpenStudy (anonymous):

The second equation is easier to work with if you multiply through by 100 to get rid of the decimals. In other words, 23x + 45y = 2436 From the first equation x = 58 - y Substitute that into the second equation and you will find your answer. Don't forget to check the answer. It ends up being a decimal.